# Dislocation-mediated short-range order evolution during thermomechanical processing

### v1

@@ Abstract
Thermomechanical processing alters the microstructure of metallic alloys through coupled plastic deformation and thermal exposure, with dislocation motion driving plasticity and microstructural evolution. Our previous work[1] showed that the same dislocation motion both creates and destroys chemical short-range order (SRO), driving alloys into far-from-equilibrium SRO states. However, the connection between this dislocation-mediated SRO evolution and processing parameters remains largely unexplored. Here, we perform large-scale atomistic simulations of thermomechanical processing of equiatomic TiTaVW to determine how temperature and strain rate control SRO via competing creation ($\Gamma$) and annihilation ($\lambda$) rates. Using machine learning interatomic potentials and information-theoretic metrics, we quantify that the magnitude and chemical character of SRO vary systematically with these parameters. We identify two regimes: a low-temperature regime with weak strain-rate sensitivity, and a high-temperature regime in which reduced dislocation density and increased screw character amplify chemical bias and accelerate SRO formation. The resulting steady-state SRO is far-from-equilibrium and cannot be produced by equilibrium thermal annealing. Together, these results provide a mechanistic and predictive link between processing parameters, dislocation physics, and SRO evolution in chemically complex alloys.
@@ 1. Introduction
Thermomechanical processing (TMP) is a widely employed manufacturing technique in which alloys are strained at controlled temperatures to induce microstructural evolution via plastic deformation[2, 3, 4, 5, 6, 7]. Plastic deformation is most often governed by dislocation motion; therefore, the mechanical response of a material undergoing TMP reflects dislocation behavior. That behavior is controlled by temperature and strain rate — the two primary parameters of TMP[8, 9, 10]. For example, at low temperatures, dislocation motion is primarily governed by the Peierls barrier, whereas at high temperatures dislocations become thermally activated and can overcome this barrier more readily, reducing the flow stress at a fixed strain rate[11]. Furthermore, increasing temperature activates qualitatively different modes of dislocation motion such as diffusion-assisted climb, and cross-slip — which are largely suppressed at low temperatures[12]. High strain rates, on the other hand, raise the flow stress so that dislocations surmount Peierls barriers with reduced thermal assistance (i.e., athermal glide). Increasing strain rate also promotes dislocation multiplication in order to sustain the high rate of plastic deformation[13]. This leads to the formation of dense dislocation forests and junctions, many of which are sessile, that hinder the motion of glissile dislocations, resulting in strain hardening.
@@
Fundamentally, dislocation motion is a bond-breaking and bond-forming process that alters the local chemical environment in the wake of moving dislocation cores. A large number of dislocations traversing through the system can thus modify the relative population of certain local chemical motifs, thereby altering spatial chemical correlations at the nanoscale — commonly referred to as chemical short-range order (SRO) [14, 15, 16]. SRO has been reported to influence several materials properties, including mechanical strength[17, 18, 19, 20, 21], radiation damage resistance[22, 23], catalysis[24], and corrosion resistance[25, 26, 27], especially in alloys with pronounced chemical complexity, such as high-entropy alloys[28]. In ref. [1], we demonstrated that dislocation motion both creates and destroys SRO, driving alloys into far-from-equilibrium SRO states. Notably, even alloys initialized as chemically random solid solutions evolve SRO under plastic deformation. The mechanism underlying this phenomenon was attributed to the chemically biased nature of dislocation motion in alloys, which promotes certain motifs over others[29, 30, 31]. As described earlier, dislocation behavior is strongly dependent on temperature and strain rate. Consequently, a natural correlation must exist between these TMP parameters and the evolution of SRO via dislocation motion. However, this relationship remains obscure in the literature, leaving a critical gap in our understanding of how SRO evolves in metallic alloys. Without a mechanistic understanding of how dislocation motion affects SRO under different TMP conditions, interpretations of deformation experiments risk conflating intrinsic mechanical response with concurrent changes in SRO, and models cannot reliably predict regimes where nonequilibrium SRO dominates. This gap limits transferability across alloys and invites equilibrium-based extrapolations[32, 33, 34] that are systematically misleading for TMP.
@@
Here, we investigate the effects of temperature and strain rate on the evolution of SRO in equiatomic TiTaVW alloy during TMP using large-scale atomistic simulations. These simulations — performed on the Frontier exascale supercomputer[35] — employ a high-fidelity machine learning interatomic potential that was trained on ab initio data tailored to capture the structure and chemistry relevant to TMP (section 2.1). We select equiatomic TiTaVW as a canonical refractory alloy with a high Peierls barrier and screw-dominated plasticity, providing a stringent testbed for chemically-biased dislocation motion. In sections 3.13.3 we demonstrate how TMP processing parameters influence SRO evolution. The dislocation motion mechanisms underlying SRO evolution are discussed in section 3.4. Finally, in section 3.5, we implement an empirical model based on our simulations results to explore steady-state SRO under conditions relevant to practical manufacturing processes.
@@ 2.1. Machine learning interatomic potential
To conduct accurate atomistic simulations of equiatomic TiTaVW alloy, we developed a machine learning interatomic potential tailored to capture the chemistry-sensitive mechanisms governing dislocation behavior under thermomechanical loading — mechanisms that conventional empirical potentials fail to represent accurately[14, 36]. The training dataset for the potential included both perfect and defect configurations, carefully curated to capture the chemical and structural complexity relevant to the TMP of TiTaVW alloy. The perfect configurations included body-centered cubic (bcc), face-centered cubic (fcc), and hexagonal close-packed (hcp) solid solutions, pure elemental bcc structures, and B2 structures representing all binary combinations. The fcc and hcp phases were included to account for phase stability, with fcc supercells specifically added due to their demonstrated ability to accurately capture screw dislocation core characteristics[37, 38].
@@
Defect configurations were chosen based on well-established mechanisms governing plastic deformation in bcc systems, which is primarily mediated by the motion of $\frac{1}{2}\langle 111\rangle$ screw dislocations. The properties of these dislocations are strongly influenced by the stacking-fault energy (SFE) landscapes on the $(110)$ and $(112)$ planes[39, 40]. Accordingly, we included configurations with stacking faults along the $[111]$ direction on these planes. However, SFE curves alone are insufficient to fully represent the highly anisotropic core structures of dislocations in bcc systems[41]. To address this, we incorporated configurations containing $\frac{1}{2}\langle 111\rangle\{110\}$ screw dislocations generated via the dislocation dipole method[42], as well as configurations with $\frac{1}{2}\langle 111\rangle\{110\}$ edge dislocation dipoles for their relevance in bcc plasticity[43]. Finally, we included bcc configurations with both vacancies and interstitials (placed at either tetrahedral or octahedral sites) to account for defect-mediated mechanisms relevant to dislocation motion.
@@
For all solid-solution phases, elemental compositions were sampled across the full TiTaVW local motif space using the motif-based algorithm of ref. [44]. To bias the training dataset toward near equiatomic compositions, the procedure was adapted by first generating a diverse set of chemically random solutions and then allowing extra-cell swaps between configurations. This allowed greater flexibility to increase motif diversity, especially in low-data regimes, while favoring near equiatomic configurations to be represented.
@@
The lattice constants of these structures were adjusted using Vegard’s law to the sampled compositions. To mimic thermal vibrations, atomic displacements were applied by randomly perturbing atoms following the methodology in ref. [44]. Isotropic lattice expansion was further applied to imitate both room-temperature and high-temperature thermal expansion[36].
@@
Density functional theory (DFT) calculations were performed to obtain energies and atomic forces for all configurations in the dataset. These calculations used the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional[45] and projector-augmented wave (PAW) pseudopotentials[46], as implemented in VASP[47, 48, 49, 50, 51]. The resulting DFT energies and forces were used to train a machine learning interatomic potential based on the atomic cluster expansion framework, with hyperparameter optimized to balance both accuracy and generalization[44, 52, 53].
@@
All molecular dynamics and Monte Carlo simulations reported here were conducted using the developed machine learning interatomic potential.
@@ 2.2. Thermomechanical processing simulations
The central simulations underpinning the results of this paper were molecular dynamics simulations of uniaxial compressive deformation on bcc single crystals of TiTaVW alloy using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS)[54]. Simulation cells were initialized by randomly populating bcc lattice sites with Ti, Ta, V, and W atoms in equiatomic proportions. The initial simulation box measured $53a_{\text{o}}\times 106a_{\text{o}}\times 212a_{\text{o}}$ along the $[100]$, $[010]$, and $[001]$ directions, respectively, where $a_{\text{o}}$ is the temperature-dependent lattice constant. Each configuration contained approximately 2.4 million atoms, with periodic boundary conditions applied in all directions.
@@
Figure 1: Thermomechanical processing of equiatomic TiTaVW alloy. a) Uniaxial compression of TiTaVW at $300\,\text{K}$ under a constant strain rate of $10^{9}\,\text{s}^{-1}$. After each unit strain, the deformation axis is rotated to align with the direction of largest dimension. Dislocation networks are shown as green lines. b) Stress-strain curve shows the onset of steady-state flow stress. c) Dislocation density also reaches steady-state with continued deformation. d) Evolution of SRO ($D_{\text{sro}}$) with strain, eventually reaching a steady-state SRO ($D^{\infty}_{\text{sro}}$). The dashed orange curve shows the fit to eq. 5, capturing the observed $D_{\text{sro}}$ evolution.
The dislocation-seeded structures were first equilibrated at constant temperatures $T$ ranging from 300 K to 2,300 K and zero pressure using a Nosé-Hoover thermostat and a Nosé-Hoover barostat, respectively, for 100 ps, allowing isotropic volume relaxation. After equilibration, mechanical deformation[55, 56] was applied in compression along the $[001]$ axis at constant strain rates $\dot{\epsilon}$ ranging from $10^{8}\,\text{s}^{-1}$ to $10^{9}\,\text{s}^{-1}$, while maintaining zero lateral pressure via the Nosé-Hoover barostat in the $[100]$ and $[010]$ directions. Temperature, which remained the same as the equilibration temperature, was controlled using a Nosé-Hoover thermostat throughout deformation. For all thermostats and barostats applied throughout the simulations, we used a damping parameter of 10 ps and a chain length of 3.
After every unit strain, the compression axis was rotated sequentially from $[001]$ to $0\bar{1}0]$, then to $[100]$, and again back to $[001]$. This procedure was repeated until a total true strain of 7 was reached. This cyclic deformation process (the first three cycles of which are illustrated in fig. 1a) mimics a “metal kneading” protocol[56]. The crystal, initially shaped with a dimension ratio of 1:2:4, is compressed along its longest axis, and the process is repeated along the new longest dimension in each subsequent cycle, preserving nearly constant volume and minimizing artifacts due to periodic boundary conditions. The integration timestep of the simulations was set to 5 fs.
Dislocations in the systems were analyzed using the Dislocation Extraction Algorithm (DXA)[57]. Prior to dislocation analysis, the atomic structures were denoised using a score-based denoising algorithm[58] for 8 steps. The cumulative edge-to-screw character ratio of dislocations in each simulation was quantified using the following expression:
@@
All visualizations were performed using the Open Visualization Tool (OVITO)[59].
@@ 2.3. Annealing simulations
Simulations were initialized with a random equiatomic distribution of atoms on the bcc lattice. Thermal equilibration was achieved through atom-swap attempts between atoms of different elements, with acceptance governed by the Metropolis criterion[60, 61] probability $\exp(-\Delta E/k_{\text{B}}T)$, where $\Delta E$ is the energy change associated with the swap, $k_{\text{B}}$ is the Boltzmann constant, and $T$ is the temperature. Each simulation was carried out for 360,000 steps (i.e., 80 swaps per atom), with the first 180,000 steps used for initial equilibration. Seven statistically uncorrelated configurations were sampled every 30,000 steps during the remaining 180,000 steps, yielding equilibrium SRO states analogous to those obtained from thermal annealing.

### v2

@@ Abstract
Thermomechanical processing alters the microstructure of metallic alloys through coupled plastic deformation and thermal exposure, with dislocation motion driving plasticity and microstructural evolution. Our previous work[islam_nonequilibrium_2025] showed that the same dislocation motion both creates and destroys chemical short-range order (SRO), driving alloys into far-from-equilibrium SRO states. However, the connection between this dislocation-mediated SRO evolution and processing parameters remains largely unexplored. Here, we perform large-scale atomistic simulations of thermomechanical processing of equiatomic TiTaVW to determine how temperature and strain rate control SRO via competing creation ($\Gamma$) and annihilation ($\lambda$) rates. The simulations employ systems containing 2.4 million atoms and utilize a machine learning interatomic potential optimized to capture chemical complexity through the motif-based sampling technique. Using information-theoretic metrics, we quantify that the magnitude and chemical character of SRO vary systematically with processing parameters. We identify two regimes: a low-temperature regime with weak strain-rate sensitivity, and a high-temperature regime in which reduced dislocation density and increased screw character amplify chemical bias and accelerate SRO formation. The resulting steady-state SRO is far-from-equilibrium and cannot be produced by equilibrium thermal annealing. Together, these results provide a mechanistic and predictive link between processing parameters, dislocation physics, and SRO evolution in chemically complex alloys.
Keywords: Thermomechanical processing; Short-range order; Dislocation structure; Machine learning interatomic potentials; Atomistic simulations
@@ 1. Introduction
Thermomechanical processing (TMP) is a widely employed manufacturing technique in which alloys are strained at controlled temperatures to induce microstructural evolution via plastic deformation[rackwitz_understanding_2025, trink_processing_2023, yamanaka_stacking-fault_2017, sun_thermomechanical_2016, schuh_mechanical_2015, shahmir_effect_2016]. Plastic deformation is most often governed by dislocation motion; therefore, the mechanical response of a material undergoing TMP reflects dislocation behavior. That behavior is controlled by temperature and strain rate — the two primary parameters of TMP[rizzardi_mild--wild_2022, kubilay_high_2021, lu_relative_2021]. For example, at low temperatures, dislocation motion is primarily governed by the Peierls barrier, whereas at high temperatures dislocations become thermally activated and can overcome this barrier more readily, reducing the flow stress at a fixed strain rate[marian_dynamic_2004]. Furthermore, increasing temperature activates qualitatively different modes of dislocation motion such as diffusion-assisted climb, and cross-slip — which are largely suppressed at low temperatures[bulatov_computer_2006]. High strain rates, on the other hand, raise the flow stress so that dislocations surmount Peierls barriers with reduced thermal assistance (i.e., athermal glide). Increasing strain rate also promotes dislocation multiplication in order to sustain the high rate of plastic deformation[fan_strain_2021]. This leads to the formation of dense dislocation forests and junctions, many of which are sessile, that hinder the motion of glissile dislocations, resulting in strain hardening.
@@
Fundamentally, dislocation motion is a bond-breaking and bond-forming process that alters the local chemical environment in the wake of moving dislocation cores. A large number of dislocations traversing through the system can thus modify the relative population of certain local chemical motifs, thereby altering spatial chemical correlations at the nanoscale — commonly referred to as chemical short-range order (SRO) [sheriff_quantifying_2024, sheriff2024chemicalmotif, zhou_atomic-scale_2022]. SRO has been reported to influence several materials properties, including mechanical strength[antillon_chemical_2021, antillon_chemical_2020, dasari_exceptional_2023, chen_simultaneously_2021, zheng_multi-scale_2023], radiation damage resistance[su_enhancing_2023, el_atwani_quinary_2023], catalysis[xie_highly_2019], and corrosion resistance[qiu_corrosion_2017, xie_percolation_2021, blades_tuning_2024], especially in alloys with pronounced chemical complexity, such as high-entropy alloys[george_high-entropy_2019]. In ref. [islam_nonequilibrium_2025], we demonstrated that dislocation motion both creates and destroys SRO, driving alloys into far-from-equilibrium SRO states. Notably, even alloys initialized as chemically random solid solutions evolve SRO under plastic deformation. The mechanism underlying this phenomenon was attributed to the chemically biased nature of dislocation motion in alloys, which promotes certain motifs over others[yin_atomistic_2021, utt_origin_2022, zhou_hidden_2021]. As described earlier, dislocation behavior is strongly dependent on temperature and strain rate. Consequently, a natural correlation must exist between these TMP parameters and the evolution of SRO via dislocation motion. However, this relationship remains obscure in the literature, leaving a critical gap in our understanding of how SRO evolves in metallic alloys. Without a mechanistic understanding of how dislocation motion affects SRO under different TMP conditions, interpretations of deformation experiments risk conflating intrinsic mechanical response with concurrent changes in SRO, and models cannot reliably predict regimes where nonequilibrium SRO dominates. This gap limits transferability across alloys and invites equilibrium-based extrapolations[moniri_three-dimensional_2023, chen_direct_2021, zhang_short-range_2020] that are systematically misleading for TMP.
@@
Here, we investigate the effects of temperature and strain rate on the evolution of SRO in equiatomic TiTaVW alloy during TMP using large-scale atomistic simulations. These simulations — performed on the Frontier exascale supercomputer[atchley_frontier_2023] — employ a high-fidelity machine learning interatomic potential that was trained on ab initio data tailored to capture the structure and chemistry relevant to TMP (section 2.1). We select equiatomic TiTaVW as a canonical refractory alloy with a high Peierls barrier and screw-dominated plasticity, providing a stringent testbed for chemically-biased dislocation motion. In sections 3.13.3 we demonstrate how TMP processing parameters influence SRO evolution. The dislocation motion mechanisms underlying SRO evolution are discussed in section 3.4. Finally, in section 3.5, we implement an empirical model based on our simulation results to explore steady-state SRO under conditions relevant to practical manufacturing processes.
@@ 2.1. Machine learning interatomic potential
To conduct accurate atomistic simulations of equiatomic TiTaVW alloy, we developed a machine learning interatomic potential tailored to capture the chemistry-sensitive mechanisms governing dislocation behavior under thermomechanical loading — mechanisms that conventional empirical potentials fail to represent accurately[sheriff_quantifying_2024, cao_capturing_2024]. The training dataset for the potential included both perfect and defect configurations, carefully curated to capture the chemical and structural complexity relevant to the TMP of TiTaVW alloy. The perfect configurations included body-centered cubic (bcc), face-centered cubic (fcc), and hexagonal close-packed (hcp) solid solutions, pure elemental bcc structures, and B2 structures representing all binary combinations. The fcc and hcp phases were included to account for phase stability, with fcc supercells specifically added due to their demonstrated ability to accurately capture screw dislocation core characteristics[Zotov_2024, wang2022tamingscrewdislocationcores].
@@
Defect configurations were chosen based on well-established mechanisms governing plastic deformation in bcc systems, which is primarily mediated by the motion of $\frac{1}{2}\langle 111\rangle$ screw dislocations. The properties of these dislocations are strongly influenced by the stacking-fault energy (SFE) landscapes on the $(110)$ and $(112)$ planes[hirth1983theory, vitek1968intrinsic]. Accordingly, we included configurations with stacking faults along the $[111]$ direction on these planes. However, SFE curves alone are insufficient to fully represent the highly anisotropic core structures of dislocations in bcc systems[freitas_machine-learning_2022]. To address this, we incorporated configurations containing $\frac{1}{2}\langle 111\rangle\{110\}$ screw dislocations generated via the dislocation dipole method[yin_ab_2020], as well as configurations with $\frac{1}{2}\langle 111\rangle\{110\}$ edge dislocation dipoles for their relevance in bcc plasticity[lee_strength_2021]. Finally, we included bcc configurations with both vacancies and interstitials (placed at either tetrahedral or octahedral sites) to account for defect-mediated mechanisms relevant to dislocation motion.
@@
For all solid-solution phases, elemental compositions were sampled across the full TiTaVW local motif space using the motif-based algorithm of ref. [sheriff_machine_2025]. To bias the training dataset toward near equiatomic compositions, the procedure was adapted by first generating a diverse set of chemically random solutions and then allowing extra-cell swaps between configurations. This allowed greater flexibility to increase motif diversity, especially in low-data regimes, while favoring near equiatomic configurations to be represented.
@@
The lattice constants of these structures were adjusted using Vegard’s law to the sampled compositions. To mimic thermal vibrations, atomic displacements were applied by randomly perturbing atoms following the methodology in ref. [sheriff_machine_2025]. Isotropic lattice expansion was further applied to imitate both room-temperature and high-temperature thermal expansion[cao_capturing_2024].
@@
Density functional theory (DFT) calculations were performed to obtain energies and atomic forces for all configurations in the dataset (see Supplementary Information table 1). These calculations used the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional[PBE] and projector-augmented wave (PAW) pseudopotentials[PAW], as implemented in VASP[VASP_1, VASP_2, VASP_3, VASP_4, VASP_PAW]. The resulting DFT energies and forces were used to train a machine learning interatomic potential based on the atomic cluster expansion framework, with hyperparameters optimized to balance both accuracy and generalization[sheriff_machine_2025, drautz_atomic_2019, lysogorskiy_performant_2021].
@@
All molecular dynamics and Monte Carlo simulations reported here were conducted using the developed machine learning interatomic potential. For more details on the training dataset structures and their construction, thermal noise and expansion augmentations, DFT simulation setup, and potential performance, we refer the reader to Supplementary Information section 1.
@@ 2.2. Thermomechanical processing simulations
The central simulations underpinning the results of this paper were molecular dynamics simulations of uniaxial compressive deformation on bcc single crystals of TiTaVW alloy using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS)[thompson_lammps_2022]. Simulation cells were initialized by randomly populating bcc lattice sites with Ti, Ta, V, and W atoms in equiatomic proportions. The initial simulation box measured $53a_{\text{o}}\times 106a_{\text{o}}\times 212a_{\text{o}}$ along the $[100]$, $[010]$, and $[001]$ directions, respectively, where $a_{\text{o}}$ is the temperature-dependent lattice constant. Each configuration contained approximately 2.4 million atoms, with periodic boundary conditions applied in all directions.
@@
Figure 1: Thermomechanical processing of equiatomic TiTaVW alloy. a) Uniaxial compression of TiTaVW at $300\,\text{K}$ under a constant strain rate of $10^{9}\,\text{s}^{-1}$. After each unit strain, the deformation axis is rotated to align with the direction of largest dimension. Dislocation networks are shown as green lines. b) Stress-strain curve shows the onset of steady-state flow stress. c) Dislocation density also reaches steady-state with continued deformation. d) Evolution of SRO ($D_{\text{sro}}$) with strain, eventually reaching a steady-state SRO ($D^{\infty}_{\text{sro}}$). The dashed orange curve shows the fit to eq. 5, capturing the observed $D_{\text{sro}}$ evolution. During fitting, $D^{\text{o}}_{\mathrm{sro}}$ was fixed at $D_{\mathrm{sro}}(\epsilon=0)$.
The dislocation-seeded structures were first equilibrated at constant temperatures $T$ ranging from 300 K to 2,300 K and zero pressure using a Nosé-Hoover thermostat and a Nosé-Hoover barostat, respectively, for 100 ps, allowing isotropic volume relaxation. After equilibration, mechanical deformation[zepeda-ruiz_atomistic_2021, zepeda-ruiz_probing_2017] was applied in compression along the $[001]$ axis at constant strain rates $\dot{\epsilon}$ ranging from $10^{8}\,\text{s}^{-1}$ to $10^{9}\,\text{s}^{-1}$, while maintaining zero lateral pressure via the Nosé-Hoover barostat in the $[100]$ and $[010]$ directions. Temperature, which remained the same as the equilibration temperature, was controlled using a Nosé-Hoover thermostat throughout deformation. For all thermostats and barostats applied throughout the simulations, we used a damping parameter of 10 ps and a chain length of 3.
After every unit strain, the compression axis was rotated sequentially from $[001]$ to $0\bar{1}0]$, then to $[100]$, and again back to $[001]$. This procedure was repeated until a total true strain of 7 was reached. This cyclic deformation process (the first three cycles of which are illustrated in fig. 1a) mimics a “metal kneading” protocol[zepeda-ruiz_probing_2017]. The crystal, initially shaped with a dimension ratio of 1:2:4, is compressed along its longest axis, and the process is repeated along the new longest dimension in each subsequent cycle, preserving nearly constant volume and minimizing artifacts due to periodic boundary conditions. A full deformation sequence is provided in Supplementary Video 1. The integration timestep of the simulations was set to 5 fs, which is $\sim$1/37th of the period of the highest-frequency phonon mode of equiatomic TiTaVW at 300K.
Dislocations in the systems were analyzed using the Dislocation Extraction Algorithm (DXA)[stukowski_extracting_2010]. Prior to dislocation analysis, the atomic structures were denoised using a score-based denoising algorithm[hsu_score-based_2024] for 8 steps. The full dislocation evolution during a simulation is shown in Supplementary Video 2. The cumulative edge-to-screw character ratio of dislocations in each simulation was quantified using the following expression:
@@
All visualizations were performed using the Open Visualization Tool (OVITO)[stukowski_visualization_2009].
@@ 2.3. Annealing simulations
Simulations were initialized with a random equiatomic distribution of atoms on the bcc lattice. Thermal equilibration was achieved through atom-swap attempts between atoms of different elements, with acceptance governed by the Metropolis criterion[metropolis_equation_1953, hastings_monte_1970] probability $\exp(-\Delta E/k_{\text{B}}T)$, where $\Delta E$ is the energy change associated with the swap, $k_{\text{B}}$ is the Boltzmann constant, and $T$ is the temperature. Each simulation was carried out for 360,000 steps (i.e., 80 swaps per atom), with the first 180,000 steps used for initial equilibration. Seven statistically uncorrelated configurations were sampled every 30,000 steps during the remaining 180,000 steps, yielding equilibrium SRO states analogous to those obtained from thermal annealing.